Can someone PLEASE answer this question? It would be very helpful, Thanks(: The question is:
Victor had a 15 kilometer start on Michael. How long will it take Michael to catch Victor if Victor is traveling at 35 kilometers per hour and Michael is traveling at 95 kilometers per hour?
Thanks again!!

Answer by jorel1380(3719)

95n=35n+15
60n=15
n=15/60 hour or 15 minutes..
Victor=15+1/4(35)=15+8.75=23.75
Michael=1/4(95)=23.75